OpenStudy (anonymous):

A fair die is rolled 4 times. What is the probability that a 4 is obtained on at least one of the rolls? Round your answer to three decimal places.

OpenStudy (anonymous):

P(4 on at least one roll)=1-P(4 on no rolls) = 1-(5/6)^4 = 0.517747 = 52%